# Crystalline Spectral Form Factors

### v1

@@ Abstract
Introduction— Spectral statistics and random matrix theory (RMT) are cornerstones of quantum chaos and quantum statistical mechanics [Porter, Mehta, Forrester, Haake, Stockmann, Guhr:1997ve, DAlessio:2015]. Most quantum systems fall into one of the four major RMT classes  [^1] A complete classification of random matrix ensembles reveals additional spectral features [Altland:1997, Kawabata:2018gjv]. distinguished by the probability density function of eigenvalue (energy level) spacings $s$, $P(s)\sim s^{\beta}$ as $s\to 0$. Quantum chaos is associated with Wigner-Dyson statistics, which corresponds to $\beta=1,2$ or $4$ and implies level repulsion [Bohigas:1984]. On the other hand, integrable systems lack level correlations and follow Poisson statistics with $\beta=0$ [Berry:1977a]. Since the inception of RMT in the framework of nuclear physics [Weidenmuller:2008vb], these four universal classes have been observed in a wide range of quantum systems and provided essential tools for quantum many-body dynamics [Srednicki:1994, Rigol:2008, Sachdev, Nandkishore:2014, Abanin:2018], quantum information theory [Collins:2016], quantum optics [Carmichael], quantum transport [Beenakker:1997], quantum gravity [Brezin:1978, Cotler:2016fpe, Saad:2019lba, Cotler:2020ugk, Belin:2023efa, Jafferis:2025vyp], and many other areas of physics.
However, some quantum systems fall outside the standard RMT classes with $\beta=0,1,2$ or $4$. One class of remarkable examples of non-conventional yet universal level statistics are provided by quantum cellular automata [Farrelly:2019zds, Arrighi:2019uor, Gopalakrishnan:2018, Gopalakrishnan:2018rfu, Gopalakrishnan:2018wyg, Alba:2019okd, Iaconis:2019hab, Iaconis:2021, Feldmeier:2020xxb, Hillberry:2024xpb] and random permutation circuits [Bertini:2024tuo, Szasz-Schagrin:2025dil, Bertini:2025vvz]. In a properly chosen basis, discrete unitary time evolution of these models breaks down into large periodic cycles, so their quasi-energy spectrum reduces to sets of roots of unity  [^2] Possibly shifted by a constant phase.. Once the length of the largest cycle is comparable to the dimension of the Hilbert space (or dynamics is restricted to one of the invariant subspaces), quasi-energy levels become equally spaced on the unit circle. The probability density function of such a “picket fence” ensemble is exactly zero in the vicinity of level spacing $s=0$, so this behavior corresponds to $\beta=\infty$. At the same time, these models are easily engineered and exhibit non-trivial dynamics. So, it becomes increasingly interesting to understand the emergence of quantum chaos in the entire range $\beta\in(0,\infty)$, and to find analytically tractable models that show behavior that interpolates between $\beta=\infty$ and standard RMT classes.
@@
Here, $U(t)=(U(1))^{t}$ is the evolution operator for integer time $t$, $E_{n}$ are the corresponding (quasi)energies, $d$ is the dimension of the Hilbert space, and $\langle\ldots\rangle$ is the ensemble average over statistically similar systems. Since the models that we study exhibit long-range (in energy) crystalline-like correlations of the eigenvalues of $U=U(1)$, we will employ the analogy with Bragg diffraction and treat the SFF as the structure factor of an eigenvalue crystal [Ashcroft] which is subject to fluctuations. Following this analogy, quasi-energies $E_{m}$ correspond to positions of atoms, and time $t$ corresponds to the momentum change of the scattered wave.
Figure 1: The SFF of the circular $\beta$-ensemble, which describes a Coulomb crystal of $d=512$ repulsive eigenvalues at inverse temperature $\beta=500$. We average over $10^{6}$ samples. Dashed line shows the damping Debye-Waller factor. Inset: deviation of the SFF from the plateau (zoom in the same graph).
The SFF is the simplest nontrivial quantity that captures eigenvalue correlations over the entire spectrum. In integrable systems the SFF decreases, in many cases monotonically, from $K(0)=d^{2}$ to $K(t)\to d$ as $t\to\infty$. In conventionally quantum chaotic systems, it exhibits a telltale “dip—ramp—plateau” structure, which reflects level repulsion [Haake, Stockmann, Guhr:1997ve]. However, none of these standard SFF types show structure at long times $t$ that are large multiples of $d$. The models that we study in this Letter all can exhibit crystalline-like oscillations of the SFF extending to $t\gg d$ due to extremely strong level repulsion (Fig. 1). Similarly to Bragg diffraction from a real crystal, these oscillations are damped due to the weak fluctuations in the level patterns once we move away the exact $\beta=\infty$ limit [Ashcroft]. We calculate the damping Debye-Waller factor for all mentioned models and confirm it with numerical results. We also extend this calculation to higher-order derivatives in the case of the circular $\beta$-ensemble and show that well-known features of the SFF for the standard cases $\beta=1,2,4$ are remnants of Bragg peaks.
Coulomb crystal of eigenvalues— The circular $\beta$-ensemble extends the Wigner-Dyson level statistics of unitary operators from $\beta=1,2,4$ to an arbitrary real non-negative $\beta$. The joint probability distribution for the quasi-energies $E_{n}$ of a unitary matrix drawn from the circular $\beta$-ensemble is described by the $d$-particle Coulomb gas confined to a unit circle at temperature $T=1/\beta$  [^3] The equilibration of the Coulomb gas is described by the Dyson Brownian motion [Porter, Mehta, Forrester] and reproduces probability distribution (2). We emphasize that this description does not rely on the specific properties of $\beta=1,2,4$.:
@@
We emphasize that, in general, the probability distribution of the random matrices that realize the circular $\beta$-ensemble is not invariant under basis transformations. Only the standard values $\beta=1,2,4$ correspond to basis-invariant ensembles  [^4] Physically sound basis transformations are described by an associative division algebra over real numbers. According to the Frobenius theorem, all such algebras are isomorphic to either real, complex, or quaternion numbers. These algebras correspond to the orthogonal ($\beta=1$), unitary ($\beta=2$), and symplectic ($\beta=4$) basis transformations, respectively.. Nonetheless, level statistics (2) with arbitrary $\beta>0$ does emerge in basis-specific settings, including a class of random-walk-like models [Killip:2004, Killip:2006, CMV]. This relationship allows us to sample the circular $\beta$-ensemble at low computational cost and study its spectral properties beyond $\beta=1,2,4$. Furthermore, it opens a way for experimental tests and many-body generalizations of our models: see Supplemental Material and Ref. [Kolganov:2024nzi] for examples of quantum circuits implementing the circular $\beta$-ensemble.
@@
The last block is diagonal and depends on a single real parameter uniformly distributed on the interval $(-\pi,\pi)$: $\Theta_{2l-1}=e^{i\phi_{2l-1}}\oplus 1$. This specifies the model: it is proven in [Killip:2004, Killip:2006] that the eigenvalues of unitaries (3) are distributed according to the circular $\beta$-ensemble for any $d=2l$ (the model for an odd $d$ is defined similarly).
@@
In particular, this formula recovers the exact results [Mehta] for $\tau=1$ ($\beta=1,2,4$) and $\tau=2$ ($\beta=4$). However, we emphasize that Gaussian approximation (9) works only for relatively small times $\tau\ll\beta$ (see Supplemental Material for details). This limitation explains the discrepancy between Eq. (12) and exact $\beta=1,2,4$ results at larger times, where the SFF in those specific cases has no singularities. Nevertheless, it leaves ample room for conspicuous crystalline-like behavior of the SFF when $\beta\gg 1$.
@@
Perturbed permutation circuit— There are other models interpolating between the standard RMT universality classes. Besides empirical Brody [Brody:1973, Brody:1981] and Berry-Robnik [Berry:1984a] distributions, explicit examples of intermediate level statistics include banded matrices [Moshe:1994gc, Mirlin:1996, Bogomolny:2010, Evers:2008], Pandey-Mehta [Pandey:1982br, Lenz:1990, Lenz:1991, Dupuis:1991], Rosenzweig-Porter [Rosenzweig:1960, Kravtsov:2015], and other toy models [Izrailev:1990elx, Prosen:1993, Prosen:1994, Prosen:1997, Prosen:2002, Prosen:2018, Chan:2018dzt, Friedman:2019gyi, Sierant:2020]. However, these models are restricted to the interval between $\beta=0$ and $\beta=2$. Hence, they cannot produce the crystalline-like behavior of the SFF that is present at large $\beta$. We suggest a new simple model, which interpolates between $\beta=2$ and $\beta=\infty$:
@@
Here, $H$ is a random $d\times d$ Hermitian matrix drawn from the Gaussian Unitary Ensemble (GUE) with probability density $\propto\exp\!\left[-(d/2)\mathrm{Tr}HH^{\dagger}\right]$, $S$ is a classical permutation of basis vectors supplemented with independent and uniformly distributed random phase rotations, $S|k\rangle=e^{i\phi_{k}}|\pi(k)\rangle$, and $0\leq g\leq\pi/2$ is a real parameter controlling the strength of this GUE perturbation to the permutation. For simplicity, we consider global matrices $H$ and $S$; however, model (13) is straightforwardly reproduced by a local shallow circuit  [^6] We discuss a straightforward local generalization of model (13) in the Supplemental Material. Besides, local permutation circuits with exponentially large cycles can be constructed using primitive polynomials over the binary field, e.g., see [Ippoliti:2025oro, Golomb].. We also assume that the permutation $\pi(k)$ is fixed but drawn at random. In this case, the typical length of its largest cycle is comparable to $d$ [Ford:2022], and the Heisenberg time $t_{H}=\mathcal{O}(d)$.
@@
Random Lax matrix ensemble— The last example of the crystalline SFF is given by the ensemble of random Lax matrices $U$ of the Ruijsenaars-Schneider model [Bogomolny:2009, Bogomolny:2011], which also emerges as quantization of interval-exchange maps [Bogomolny:2004, Giraud:2004]:
@@
where $p_{m}$ are independent random variables uniformly distributed between $\pi$ and $\pi$, $d$ is the dimension of the Hilbert space, and parameter $0<g<1$ controls level repulsion. In the limit $g\to 0$, unitary (15) becomes diagonal and has Poisson level statistics. In the limit $g\to 1$, it turns into a cyclic shift operator supplemented with random phase rotations: $U_{mn}=e^{ip_{m}}\delta_{m+1,n}$, so has a zero-temperature eigenvalue crystal. Thus these two limits correspond to $\beta=0$ and $\beta=\infty$, respectively. For intermediate values of $g$, the spectral statistics of $U$ is described by a gas of hard rods [Bogomolny:2011]:
@@
This result is exact in the limit $d\to\infty$ [Bogomolny:2011] and agrees with numerical results for $d\gg 1$, e.g., see Fig. 4. Similarly to the SFF of the Coulomb gas (9), this SFF oscillates in an earlier time regime with period $t_{H}\approx d$ when $1-g\ll 1$. Not so similarly, $K(t)/d$ and all of its time derivatives remain finite for $0<g<1$ and any $\tau=t/d>0$ in the limit $d\to\infty$. For $0<t\ll d/(1-g)$ the peaks in $K(t)/d$ are of height $\sim[d/((1-g)t)]^{2}\gg 1$ and thus strong. For later times, $K(t)/d$ continues to weakly oscillate, but now with a longer period $d/g$ and amplitude $\sim d/((1-g)t)\ll 1$.
@@
Conclusion— We have presented three models of “crystalline” level statistics associated with oscillations of the SFF. For all three models, we have analytically derived expressions for the Debye-Waller factor, which suppresses oscillations of the SFF, and the new time scale $t_{*}\gg t_{H}$, at which the SFF approaches the plateau. The key ingredient ensuring such a crystalline-like behavior is the strong level repulsion. Our examples indicate that this behavior occurs in some broader class of chaotic quantum systems. We also emphasize that this class is not restricted to unitary discrete-time evolution, although the examples we gave are of this type. In particular, the necessary requirement for crystalline level statistics is satisfied by the Gaussian $\beta$-ensemble at large $\beta$ [Dumitriu:2002], the weakly-disordered one-dimensional Anderson localization model [Torres-Herrera:2019], an ensemble of random Lax matrices of Calogero-Moser model [Bogomolny:2009, Bogomolny:2011], and inhomogeneous discrete-time quantum walks [Cantero:2010, Linden:2009, Ambarish:2017, Gong:2022, Kempe:2003, Venegas-Andraca:2012zkr, Aharonov:1993gdz, Ambainis:2001, Ryan:2005].
We expect that our work paves a way for further investigations in various directions. First, crystalline-like eigenvalue statistics implies that dephasing in the eigenstate basis is delayed until time $t_{*}\gg t_{H}$, so it suggests a careful study of the behavior of correlation functions, entanglement, complexity, and other essential tools of quantum dynamics, for our examples between $t_{H}$ and $t_{*}$. Second, it would be interesting to find and investigate many-body systems with crystalline-like level statistics intermediate between $\beta\sim 1$ and $\beta=\infty$. A straightforward generalization of the discrete-walk-like model (3) to the case of free non-interacting fermions, which we describe in the Supplemental Meterial, resembles other toy models of the many-body SFF [Liao:2020lac, Winer:2020mdc, Flynn:2024thf, Ikeda:2024mgb] and serves as a good starting point for this endeavor. Besides, models of intermediate level statistics between $\beta=0$ and $\beta\sim 1$ play an important role for the studies of many-body-localization transitions [Oganesyan:2006, Atas:2012, Buijsman:2019, Prakash:2020psj]. It would be thrilling to find an analog of this transition between $\beta\sim 1$ and $\beta=\infty$. Finally, recent experimental advancements [Dong:2024yaf] make it tempting to experimentally confirm the crystalline-like behavior of the SFF in our models.
Acknowledgements— The authors thank Sarang Gopalakrishnan and Amit Vikram for helpful discussions. D.A.H. was supported in part by NSF QLCI grant OMA-2120757.

### v2

@@ Abstract
Introduction— Spectral statistics and random matrix theory (RMT) are cornerstones of quantum chaos and quantum statistical mechanics [1, 2, 3, 4, 5, 6, 7]. Most quantum systems fall into one of the four major RMT classes  [^1] A complete classification of random matrix ensembles reveals additional spectral features [98, 99]. distinguished by the probability density function of eigenvalue (energy level) spacings $s$, $P(s)\sim s^{\beta}$ as $s\to 0$. Quantum chaos is associated with Wigner-Dyson statistics, which corresponds to $\beta=1,2$ or $4$ and implies level repulsion [9]. On the other hand, integrable systems and localized systems lack level correlations and follow Poisson statistics with $\beta=0$ [10]. Since the inception of RMT in the framework of nuclear physics [11], these four universal classes have been observed in a wide range of quantum systems and provided essential tools for quantum many-body dynamics [12, 13, 14, 15, 16], quantum information theory [17], quantum optics [18], quantum transport [19], quantum gravity [20, 21, 22, 23, 24, 25], and many other areas of physics.
However, some quantum systems fall outside the standard RMT classes with $\beta=0,1,2$ or $4$. One class of remarkable examples of non-conventional yet universal level statistics are provided by quantum cellular automata [26, 27, 28, 29, 30, 31, 32, 33, 34, 35] and random permutation circuits [36, 37, 38]. In a properly chosen basis, discrete-time unitary evolution of these models breaks down into large periodic cycles, so their quasi-energy spectrum reduces to sets of roots of unity  [^2] Possibly shifted by a constant phase.. Once the length of the largest cycle is comparable to the dimension of the Hilbert space (or dynamics is restricted to one of the invariant subspaces), quasi-energy levels become equally spaced on the unit circle. The probability density function of such a “picket fence” ensemble is exactly zero in the vicinity of level spacing $s=0$, so this behavior corresponds to $\beta=\infty$. At the same time, these models are readily engineered and exhibit non-trivial dynamics. So, it becomes increasingly interesting to understand the emergence of quantum chaos in the entire range $\beta\in(0,\infty)$, and also to find analytically tractable models that show behavior that interpolates between $\beta=\infty$ and standard RMT classes.
Figure 1: The SFF of the circular $\beta$-ensemble, which describes a Coulomb crystal of $d$ repulsive eigenvalues at inverse temperature $\beta$. We illustrate this for $d=512$, $\beta=500$, averaged over $10^{6}$ samples. Dashed line shows the damping Debye-Waller factor. Inset: deviation of the SFF from the plateau (zoom in the same graph).
@@
Here, $U(t)=(U(1))^{t}$ is the evolution operator for integer time $t$, $E_{n}$ are the corresponding (quasi)energies, $d$ is the dimension of the Hilbert space, and $\langle\ldots\rangle$ is the ensemble average over statistically similar systems. Since the models that we study exhibit long-range (in energy) crystalline-like correlations of the eigenvalues of $U=U(1)$, we will employ the analogy with Bragg diffraction and treat the SFF as the structure factor of an eigenvalue crystal which is subject to fluctuations  [^3] Note also an analogy between the eigenvalues of a random matrix ensemble and a system of spinless fermions with long-range repulsive potential [3], which also implies oscillations of the SFF [100, 101].. Following this analogy, quasi-energies $E_{m}$ correspond to positions of atoms, and time $t$ corresponds to the momentum change of the scattered wave.
The SFF is the simplest nontrivial quantity that captures eigenvalue correlations over the entire spectrum. In integrable systems the SFF decreases, in many cases monotonically, from $K(0)=d^{2}$ to $K(t)\to d$ as $t\to\infty$. In conventionally quantum chaotic systems, it exhibits a telltale “dip—ramp—plateau” structure, which reflects level repulsion [4, 5, 6]. However, none of these standard SFF types show structure at long times $t$ that are large multiples of $d$. The models that we study in this Letter all can exhibit crystalline-like oscillations of the SFF extending to $t\gg d$ due to extremely strong level repulsion (Fig. 1). Similarly to Bragg diffraction from a real crystal, these oscillations are damped due to the weak fluctuations in the level patterns once we move away the exact $\beta=\infty$ limit [41]. We calculate the damping Debye-Waller factor for all mentioned models and confirm it with numerical results. We also extend this calculation to higher-order derivatives in the case of the circular $\beta$-ensemble and show that well-known features of the SFF for the standard cases $\beta=1,2,4$ are remnants of Bragg peaks.
Coulomb crystal of eigenvalues— The circular $\beta$-ensemble extends the Wigner-Dyson level statistics of unitary operators from $\beta=1,2,4$ to an arbitrary real non-negative $\beta$. The joint probability distribution for the quasi-energies $E_{n}$ of a unitary matrix drawn from the circular $\beta$-ensemble is described by the $d$-particle Coulomb gas confined to a unit circle at temperature $T=1/\beta$  [^4] The equilibration of the Coulomb gas is described by the Dyson Brownian motion [1, 2, 3] and reproduces probability distribution (2). We emphasize that this description does not rely on the specific properties of $\beta=1,2,4$.:
@@
We emphasize that, in general, the probability distribution of the random matrices that realize the circular $\beta$-ensemble is not invariant under basis transformations. Only the standard values $\beta=1,2,4$ correspond to basis-invariant ensembles  [^5] Physically sound basis transformations are described by an associative division algebra over real numbers. According to the Frobenius theorem, all such algebras are isomorphic to either real, complex, or quaternion numbers. These algebras correspond to the orthogonal ($\beta=1$), unitary ($\beta=2$), and symplectic ($\beta=4$) basis transformations, respectively.. Nonetheless, level statistics (2) with arbitrary $\beta>0$ does emerge in basis-specific settings, including a class of random-walk-like models [44, 45, 46]. This relationship allows us to sample the circular $\beta$-ensemble at low computational cost and study its spectral properties beyond $\beta=1,2,4$. Furthermore, it opens a way for experimental tests and many-body generalizations of our models: see Supplemental Material and Ref. [47] for examples of quantum circuits implementing the circular $\beta$-ensemble.
@@
The last block is diagonal and depends on a single real parameter uniformly distributed on the interval $(-\pi,\pi)$: $\Theta_{2l-1}=e^{i\phi_{2l-1}}\oplus 1$. This specifies the model: it is proven in [44, 45] that the eigenvalues of unitaries (3) are distributed according to the circular $\beta$-ensemble for any $d=2l$ (the model for an odd $d$ is defined similarly).
@@
In particular, this formula recovers the exact results [2, 3] for $\tau=1$ ($\beta=1,2,4$) and $\tau=2$ ($\beta=4$). However, we emphasize that Gaussian approximation (9) works only for relatively small times $\tau\ll\beta$ (see Supplemental Material for details). So, in general, the estimate (12) is guaranteed to be valid only for $\tau\ll\beta$, even though properties of a given $\beta$ can significantly extend this region  [^7] For a rational $\beta/2=p/q$ with $p$ and $q$ mutually prime, the asymptotic expansion of the density-density correlation function of the circular $\beta$-ensemble, Eq. (13.226) in [3], implies that the SFF has no singularities for $\tau>p$, whereas Eq. (12) is valid for $\tau\leq p$.. This limitation explains the discrepancy between Eq. (12) and exact $\beta=1,2,4$ results at larger times, where the SFF in those specific cases has no singularities. Nevertheless, it leaves ample room for conspicuous crystalline-like behavior of the SFF for any $\beta\gg 1$.
@@
Perturbed permutation circuit— There are other models interpolating between the standard RMT universality classes. Besides empirical Brody [50, 51] and Berry-Robnik [52] distributions, explicit examples of intermediate level statistics include banded matrices [53, 54, 55, 56], Pandey-Mehta [57, 58, 59, 60], Rosenzweig-Porter [61, 62], and other toy models [63, 64, 65, 66, 67, 68, 69, 70, 71]. However, these models are restricted to the interval between $\beta=0$ and $\beta=2$. Hence, they cannot produce the crystalline-like behavior of the SFF that is present at large $\beta$. We suggest a new simple model, which interpolates between $\beta=2$ and $\beta=\infty$:
@@
Here, $H$ is a random $d\times d$ Hermitian matrix drawn from the Gaussian Unitary Ensemble (GUE) with probability density $\propto\exp\!\left[-(d/2)\mathrm{Tr}HH^{\dagger}\right]$, $S$ is a classical permutation of basis vectors supplemented with independent and uniformly distributed random phase rotations, $S|k\rangle=e^{i\phi_{k}}|\pi(k)\rangle$, and $0\leq g\leq\pi/2$ is a real parameter controlling the strength of this GUE perturbation to the permutation. For simplicity, we consider global matrices $H$ and $S$; however, model (13) is straightforwardly reproduced by a local shallow circuit  [^8] We discuss a straightforward local generalization of model (13) in the Supplemental Material. Besides, local permutation circuits with exponentially large cycles can be constructed using primitive polynomials over the binary field, e.g., see [102, 103].. We also assume that the permutation $\pi(k)$ is fixed but drawn at random. In this case, the typical length of its largest cycle is comparable to $d$ [73], and the Heisenberg time $t_{H}=\mathcal{O}(d)$.
@@
Random Lax matrix ensemble— The last example of the crystalline SFF is given by the ensemble of random Lax matrices $U$ of the Ruijsenaars-Schneider model [74, 75], which also emerges as quantization of interval-exchange maps [76, 77]:
@@
where $p_{m}$ are independent random variables uniformly distributed between $\pi$ and $\pi$, $d$ is the dimension of the Hilbert space, and parameter $0<g<1$ controls level repulsion. In the limit $g\to 0$, unitary (15) becomes diagonal and has Poisson level statistics. In the limit $g\to 1$, it turns into a cyclic shift operator supplemented with random phase rotations: $U_{mn}=e^{ip_{m}}\delta_{m+1,n}$, so has a zero-temperature eigenvalue crystal. Thus these two limits correspond to $\beta=0$ and $\beta=\infty$, respectively. For intermediate values of $g$, the spectral statistics of $U$ is described by a gas of hard rods [75]:
@@
This result is exact in the limit $d\to\infty$ [75] and agrees with numerical results for $d\gg 1$, e.g., see Fig. 4. Similarly to the SFF of the Coulomb gas (9), this SFF oscillates in an earlier time regime with period $t_{H}\approx d$ when $1-g\ll 1$. Not so similarly, $K(t)/d$ and all of its time derivatives remain finite for $0<g<1$ and any $\tau=t/d>0$ in the limit $d\to\infty$. For $0<t\ll d/(1-g)$ the peaks in $K(t)/d$ are of height $\sim[d/((1-g)t)]^{2}\gg 1$ and thus strong. For later times, $K(t)/d$ continues to weakly oscillate, but now with a longer period $d/g$ and amplitude $\sim d/((1-g)t)\ll 1$.
@@
Conclusion— We have presented three models of “crystalline” level statistics associated with oscillations of the SFF. For all three models, we have analytically derived expressions for the Debye-Waller factor, which suppresses oscillations of the SFF, and the new time scale $t_{*}\gg t_{H}$, at which the SFF approaches the plateau. The key ingredient ensuring such a crystalline-like behavior is the strong level repulsion. Our examples indicate that this behavior occurs in some broader class of chaotic quantum systems. We also emphasize that this class is not restricted to unitary discrete-time evolution, although the examples we gave are of this type. In particular, the necessary requirement for crystalline level statistics is satisfied by the Gaussian $\beta$-ensemble at large $\beta$ [78], the weakly-disordered one-dimensional Anderson localization model [79], an ensemble of random Lax matrices of Calogero-Moser model [74, 75], and inhomogeneous discrete-time quantum walks [80, 81, 82, 83, 84, 85, 86, 87, 88].
We expect that our work paves a way for further investigations in various directions. First, crystalline-like eigenvalue statistics implies that dephasing in the eigenstate basis is delayed until time $t_{*}\gg t_{H}$, so it suggests a careful study of the behavior of correlation functions, entanglement, complexity, and other essential tools of quantum dynamics, for our examples between $t_{H}$ and $t_{*}$. Second, it would be interesting to find and investigate many-body systems with crystalline-like level statistics intermediate between $\beta\sim 1$ and $\beta=\infty$. A straightforward generalization of the discrete-walk-like model (3) to the case of free non-interacting fermions, which we describe in the Supplemental Meterial, resembles other toy models of the many-body SFF [89, 90, 91, 92] and serves as a good starting point for this endeavor. Besides, models of intermediate level statistics between $\beta=0$ and $\beta\sim 1$ play an important role for the studies of many-body-localization transitions [93, 94, 95, 96]. It would be interesting to find an analog of this transition between $\beta\sim 1$ and $\beta=\infty$. Finally, recent experimental advancements [97] make it tempting to experimentally confirm the crystalline-like behavior of the SFF in our models.
Acknowledgements— The authors thank Sarang Gopalakrishnan, Amit Vikram, and Peter Forrester for helpful discussions. D.A.H. was supported in part by NSF QLCI grant OMA-2120757.